trains a and b are traveling the same direction on parallel tracks. train a is traveling at 100mph and train b is traveling at 120mph. train a passes a station at 10:25am.if train b passes the same station at 10:40am, what time will train b catch up to train a?
Answer by stanbon(48568)

trains a and b are traveling the same direction on parallel tracks.
train a is traveling at 100mph and train b is traveling at 120mph.
train a passes a station at 10:25am.if train b passes the same station at 10:40am,
what time will train b catch up to train a?
------------
Train A DATA:
Rate = 100 mph ; time = x hrs ; distance = rt = 100x miles
---------------------
Train B DATA:
Rate = 120 mph ; time = x-(1/4) hr ; distance = rt = 120x - 30 miles
-----------------------
EQUATION:
distance A = distance B
100x = 120x - 30
20x = 30
x = 3/2 hrs
---------------
B catches A at 10:25 + 1 1/2 hrs = 11:55
=======================
Cheers,
Stan H.